40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12568     Entry time: Tue Oct 18 18:56:57 2016     In reply to: 12566     Reply to this: 12571
Author: gautam 
Type: Update 
Category: General 
Subject: OM5 rotated to bypass OMC, green scatter is from window to PSL table 

[ericq, lydia, gautam]

  • We started today by checking leveling of ITMY table, all was okay on that front after the adjustment done yesterday. Before closing up, we will have detailed pictures of the current in vacuum layout
  • We then checked centering on OMs 1 and 2 (after having dither aligned the arms), nothing had drifted significantly from yesterday and we are still well centered on both these OMs
  • We then moved to the BS/PRM chamber and checked the leveling, even though nothing was touched on this table. Like in the OMC chamber, it is difficult to check the leveling here because of layout constraints, but I verified that the table was pretty close to being level using the small (clean) spirit level in two perpendicular directions
  • Beam centering was checked on OMs 3 and 4 and verified to be okay. Clearance of beam from OM4 towards the OMC chamber was checked at two potential clipping points - near the green steering mirror and near tip-tilt 2. Clearance at both locations was deemed satisfactory so we moved onto the OMC chamber
  • We decided to go ahead and rotate OM5 to send the beam directly to OM6 and bypass the partially transmissive mirror meant to send part of the AS beam to the OMC
  • In order to accommodate the new path, I had to remove a razor beam dump on the OMC setup, and translate OM5 back a little (see Attachment #1), but we have tried to maintain ~45 degree AOI on both OMs 5 and 6
  • Beam was centered on OM6 by adjusting the position of OM5. We initially fiddled around with the pitch and yaw knobs of OM4 to try and center the beam on OM5, but it was decided that it was better just to move OM5 rather than mess around on the BS/PRM chamber and introduce potential additional scatter/clipping
  • OMC table leveling was checked and verified to not have been significantly affected by todays work
  • It was necessary to loosen the fork and rotate OM6 to extract the AS beam from the vacuum chambers onto the AP table
  • AS beam is now on the camera, and looks nice and round, no evidence of any clipping. Some centering on in air lenses and mirrors on the AP table remains to be done. We are now pretty well centered on all 6 OMs and should have more power at the AS port given that we are now getting light previously routed to the OMC out as well. A quantitative measure of how much more light we have now will have to be done after pumping down and turning the PSL power back up
  • I didn't see any evidence of back-scattered light from the window even though there were hints of this previously (sadly the same can't be said about the green). I will check once again tomorrow, but this doesn't look like a major problem at the moment

Lydia and I investigated the extra green beam situation. Here are our findings.

  1. There appears to be 3 ghost beams in addition to the main beam. These ghosts appeared when we locked the X green and Y green individually, which lead us to conclude that whatever is causing this behaviour is located downstream of the periscope on the BS/PRM chamber
    Link to greenGhosts.JPG
  2. I then went into the BS/PRM chamber and investigated the spot on the lower periscope mirror. It isn't perfectly centered, but it isn't close to clipping on any edge, and the beam leaving the upper mirror on the periscope looks clean as well (only the X-arm green was used for this, and subsequent checks). The periscope mirror looks a bit dusty and scatters rather a lot which isn't ideal...
    Link to IMG_3322.JPG
  3. There are two steering mirrors on the IMC table which we do not have access to this vent. But I looked at the beam coming into the OMC chamber and it looks fine, no ghosts are visible when letting the main beam pass through a hole in one of our large clean IR viewing cards - and the angular separation of these ghosts seen on the PSL table suggests that we would see these ghosts if they exist prior to the OMC chamber on the card...
  4. The beam hits the final steering mirror which sends it out onto the PSL table on the OMC chamber cleanly - the spot leaving the mirror looks clean. However, there are two reflections from the two surfaces of the window that come back into the OMC chamber. Space constraints did not permit me to check what surfaces these scatter off and make it back out to the PSL table as ghosts, but this can be checked again tomorrow.
    Link to IMG_3326.JPG

I can't think of an easy fix for this - the layout on the OMC chamber is pretty crowded, and potential places to install a beam dump are close to the AS and IMC REFL beam paths (see Attachment #1). Perhaps Steve can suggest the best, least invasive way to do this. I will also try and nail down more accurately the origin of these spots tomorrow.


Light doors are back on for the night. I re-ran the dithers, and centered the oplevs for all the test-masses + BS. I am leaving the PSL shutter closed for the night

 

Attachment 1: OMCchamber.pdf  214 kB  Uploaded Tue Oct 18 23:45:59 2016  | Hide | Hide all
OMCchamber.pdf
Attachment 2: greenGhosts.JPG  6.057 MB  Uploaded Tue Oct 18 23:52:04 2016  | Hide | Hide all
greenGhosts.JPG
Attachment 3: IMG_3322.JPG  4.787 MB  Uploaded Tue Oct 18 23:54:49 2016  | Hide | Hide all
IMG_3322.JPG
Attachment 4: IMG_3326.JPG  5.212 MB  Uploaded Wed Oct 19 00:03:27 2016  | Hide | Hide all
IMG_3326.JPG
ELOG V3.1.3-